# Enhancement of adhesion strength through microvibrations: modeling and experiments

### v1

@@ 4.2 PDMS sample preparation
The PDMS samples utilized in the adhesion experiments were fabricated using the commercial elastomer Dow Sylgard 184®. The base:curing agent volume ratio was set to 10:1. The mixture was degassed under vacuum in a desiccator for 15 minutes at room temperature. The resin was subsequently poured into a petri dish containing a glass slide, previously activated by air plasma to favor intimate bonding without the use of any glue or intermediate layer. Finally, the resin was cured for an initial 18 h period at room temperature, followed by a second 90 min period at 75°C. The final dimensions of the PDMS sample were 2.5 x 2.5 cm in plane and 3 mm thickness.
@@ 4.3 Independent characterization of the interface rate-dependent adhesion
where $F_{H}={4E_{0}^{\ast}a^{3}}/(3R)$ is the Hertzian load, $F_{c}$ is the normal contact force as measured by the load cell and $a$ is the measured contact radius. By fitting the experimental data with the empirical Gent-Shultz power law model (Eq. 7) we found the reference speed to be $v_{0}=0.0445\;$ mm/s and $n=0.531$, as shown in Fig. 5. The inset in Fig. 5 shows the quantity $(\Delta\gamma_{eff}/\Delta\gamma_{0}-1)$ in log-log scale to highlight the power law scaling of the effective surface energy with respect to the crack speed.
In summary, the dimensional parameters of the system are reported in Table 1. Notice that the natural frequency $\omega_{n}$ and the damping ratio $\zeta$ of the free oscillator (not in contact) were obtained performing a free vibration test of the indenter. In Table 1, the effective contact modulus $E^{*}_{eff}=4.5$ MPa is reported that is a factor $\simeq 1.6$ larger then the quasi-static value determined by the JKR contact test (see A). This accounts for the fact the the substrate gets stiffer when excited in the frequency range $f\in\left[200,400\right]$ Hz and agrees well with our vibrating experiments (see B) and with independent DMA characterization of the PDMS material in Maghami et al. (2024b).
@@
Gent-Shultz exponent $n$ 0.531
@@ 4.4 Dynamic adhesion tests
Figure 6: Comparison with experiments a) Average spring reaction force ($\overline{F}_{k}$) vs. average contact radius ($\overline{a}$) curves for the dynamic tests performed with $f=300$ Hz and $X_{b}=[20,40,60,80,100]$ $\mu$m. b) Experimental pull-off force (3 test were performed per each value of amplitude $X_{b}$: data points correspond to the mean, the error bars indicate the standard deviation) vs. numerical prediction for different values of frequency ($f=[200,300,400]$ Hz). c) Force ($\overline{F}_{k}$) vs. time (t) and contact radius (a) vs. time (t) curves for the case ($X_{b}=80$ $\mu$m, $f=300$ Hz). Experiments and simulations are matched at the time instant when unloading starts ($t\simeq 120$ s). Note that each marker shape always corresponds to the same value of $X_{b}$, in all the panels. The extra orange square and right-oriented triangle in panel b) refer to $X_{b}=120$ $\mu$m and $X_{b}=160$ $\mu$m, respectively
@@
The behavior at pull-off was further scrutinized in Fig. 7a, where the pull-off force is shown as a function of the vibration amplitude for $f=300$ Hz. The black solid line refers to the model predictions while the markers (each corresponding to a certain value of $X_{b}$, consistently with Fig. 6) refer to the experimental data (the error-bar indicates one standard deviation over three independent measurements). Figure 7 b-i show micrographs of the contact area immediately prior to detachment, with the contact periphery indicated by a dashed orange circle. Pull-off is preceded by a significant reduction of contact radius in the quasi-static case and in the small amplitudes cases. As $X_{b}$ increases, the detachment takes place in a more ”abrupt” manner. The latter case, referring to $X_{b}=[100,120,160]$ $\mu$m, seems related to the saturation of the pull-off force that the model is unable to predict. It appears that above a certain vibration amplitude, $X_{b}\simeq 80$ $\mu$m for the case considered in Fig. 7a, the condition from detachment shifts from an instability point to a different mechanism. Further scrutiny of the micrographs at high vibration amplitudes reveals the presence of radial wrinkles (see close-up view for the case $X_{b}=160$ $\mu$m in Fig. 7j) close to the contact periphery. A possible interpretation is that these wrinkles are the result of local elastic instabilities on the surface due to the hoop compressive stress $\sigma_{\theta}$ generated during the unloading phase, which for $\nu=0.5$, are equal to the Hertzian component of the pressure distribution acting normal to the substrate interface ($\sigma_{\theta}=-p_{0}\sqrt{1-(r/a)^{2}}$, being $p_{0}$ the maximum stress in the Hertzian solution, $r$ the radial coordinate, see Eq. (4.130) in Maugis (2013)). Nevertheless, local instabilities, resulting in wrinkles and cavitation phenomena, have been studied theoretically and numerically in spherical adhesive contact. He et al. (2022) observed them for quasi-static loading of very confined elastic systems where $a/h\gtrsim 4$, being $h$ the thickness of the layer (while in our case we have $a/h\simeq 1)$. However, our experimental set-up involves a viscoelastic material with base excitation, therefore definitive conclusions on the origin of the observed wrinkles cannot be drawn at this point.
@@
By considering both the pull-off force and the contact area (for $X_{b}>80$ $\mu$m, we considered only the patch within the white dashed circle to be in intimate contact), we estimated the average stress at pull-off, $\overline{\sigma}_{po}$, which is plotted as a function of the vibration amplitude in the inset of Fig. 7a. The results show that $\overline{\sigma}_{po}$ increases with $X_{b}$, peaks at $X_{b}=80$ $\mu$m, and decreases as $X_{b}$ increases further. This has two implications: firstly, there is an ”optimal” vibration amplitude that maximizes the adhesive capabilities of the interface; secondly, the pull-off force saturation cannot be ascribed to the limited adhesive strength of the interface (a ”strength-limited” mechanism, see Papangelo and Ciavarella (2023)). This, together with the observation of radial wrinkles, raises the question of what is the mechanism that triggers the saturation of the pull-off force at large amplitudes. This matter goes beyond the scope of the present work, and will be subject of further investigations.
@@
Figure 7: a) Pull-off force as a function of the vibrations amplitude $X_{b}$ (each marker corresponds to a certain value of $X_{b}$, consistently with Fig. 6). The model prediction (solid line) is compared against the experimental data points. b-i) Micrographs of the contact area just prior to pull-off for each test: b) quasi-static; c) $X_{b}=20$ $\mu$m; d) $X_{b}=40$ $\mu$m; e)$X_{b}=60$ $\mu$m; f) $X_{b}=80$ $\mu$m; g) $X_{b}=100$ $\mu$m; h) $X_{b}=120$ $\mu$m; i) $X_{b}=160$ $\mu$m; j) close-up view of the radial wrinkles leading to abrupt pull-off ($X_{b}=160$ $\mu$m). The frequency was set to 300 Hz. The contrast was enhanced to allow for a better visualization
@@ 5 Discussion
In contrast, here we started with an independent characterization of the rate-dependent adhesion law under quasi-static conditions (see Fig. 5), i.e., without vibrations. By providing all the dimensional characteristic parameters of our experimental setup as input to the model, it successfully predicted the experimental results, including the time-dependent evolution of the contact radius, spring reaction force and the pull-off force, without any adjustable parameter. Nevertheless, in experiments, the pull-off force saturated above an excitation amplitude of about $100$ $\mu$m. This saturation, which could not be predicted by the model, may be related to the onset of local surface instabilities detected in the contact area snapshot at the pull-off point (see Fig. 7) perhaps due to local circumferential compressive stresses arising within the contact area. Notice that, once the indenter detached from the substrate, we could not identify any permanent damage on the PDMS surface. In order to explain the saturation of the pull-off force, Shui et al. (2020) have proposed a truncation of the Gent and Schultz rate-dependent adhesion law at a critical crack speed $v_{C}$. This scenario does not find an experimental evidence in our measurements, where the effective work of adhesion was found to monotonically increase with the crack speed (see Fig. 5). Moreover, the pull-off saturation we found appeared to be frequency dependent (see Fig. 6b), which does not support the existence of a speed cut-off in the Gent and Schultz law.
@@ 6 Conclusions
To quantitatively compare with the model predictions, we conducted an independent assessment of the rate-dependent adhesion performance of the glass sphere-PDMS substrate interface. The rigid sphere was unloaded from the PDMS substrate at several unloading rates, and the relation between the effective surface energy, $\Delta\gamma_{eff}$, and the crack speed was obtained. During the characterization, the vibration excitation of the substrate was switched off. Having obtained the parameters of the Gent and Schultz adhesion law ($v_{0}=44.5$ $\mu$m/s and $n=0.531$) for the interface, we exploited the mechanical model to predict the experimental results. This resulted in a very satisfactorily comparison, not only for single-value predictions such as the pull-off force, but also for the entire evolution of average contact radius and contact force over time.
@@ Acknowledgment
All authors were partly supported by the Italian Ministry of University and Research under the Programme “Department of Excellence” Legge 232/2016 (Grant No. CUP - D93C23000100001). A.P. and M.T. were supported by the European Union (ERC-2021-STG, “Towards Future Interfaces With Tuneable Adhesion By Dynamic Excitation” - SURFACE, Project ID: 101039198, CUP: D95F22000430006). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.C. was partly supported by the European Union through the program – Next Generation EU (PRIN-2022-PNRR, ”Fighting blindness with two photon polymerization of wet adhesive, biomimetic scaffolds for neurosensory REtina-retinal Pigment epitheliAl Interface Regeneration” - REPAIR, Project ID: P2022TTZZF, CUP: D53D23018570001).
@@ Appendix A Characterisation of the PDMS sample
Figure 8: Average spring reaction force ($\overline{F}_{k}$) vs. average contact radius ($\overline{a}$) curves for the quasi-static test (in gray) and dynamic tests performed with $f=300Hz$ and $X_{b}=[20,40,60,80,100]\mu$m and relative JKR fits (black dashed lines).
Figure 9: Evolution of the effective contact modulus $E^{*}_{eff}$ and effective work of adhesion $\Delta\gamma_{eff}$ as functions of the vibration amplitude $X_{b}$ ($f=300$ Hz). The data points were obtained by fitting the experimental force vs. contact radius curves with a JKR model, as shown in Fig. 8
Appendix B Comparison with Shui et al.
Shui et al. (2020) has proposed a slightly different model for the system in Fig. LABEL:fig1 starting from the hypothesis that the effect of the vibrating substrate can be taken into account by a forcing term $f(t)\simeq-m\omega^{2}X_{b}\sin\left(\omega t\right)$. By writing the indentation as a function of the oscillator and base position
@@
Writing the contact force $F_{c}=\frac{Ka}{2}\left(\frac{a^{2}}{R}-3\delta\right)$ and considering $k\rightarrow 0$ and $c\ll m\omega$, one obtains
@@
Nevertheless, the predictions of Eq. (25) compared very poorly with our experimental data and we were not able to clarify how Eq. (24) relates to the parameters of the Gent and Schultz adhesion law.
@@
being $\omega_{nc}=\sqrt{\left(\left.\frac{dF_{c}}{d\delta}\right|_{\delta=\overline{\delta}}\right)/m}$ which gives the ”resonance” frequency of the system. Notice that the latter is $\overline{\delta}$ dependent.

### v2

@@ 4.2 PDMS sample preparation
The PDMS samples utilized in the adhesion experiments were fabricated using the commercial elastomer Dow Sylgard 184®. The base:curing agent volume ratio was set to 10:1. The mixture was degassed under vacuum in a desiccator for 15 minutes at room temperature. The resin was subsequently poured into a petri dish containing a glass slide, previously activated by air plasma to favor intimate bonding without the use of any glue or intermediate layer. Finally, the resin was cured for an initial 18 h period at room temperature, followed by a second 90 min period at 75°C. Such a two-step protocol was adapted from Delplanque et al. (2022) and allows to reduce the shrinkage-induced deformation of the PMDS slab upon curing, preserving a flat surface. The final dimensions of the PDMS sample were 2.5 x 2.5 cm in plane and 3 mm thickness.
@@ 4.3 Independent characterization of the interface rate-dependent adhesion
where $F_{H}={4E_{0}^{\ast}a^{3}}/(3R)$ is the Hertzian load, $F_{c}$ is the normal contact force as measured by the load cell and $a$ is the measured contact radius. By fitting the experimental data with the empirical Gent-Shultz power law model (Eq. 7) we found the reference speed to be $v_{0}=0.0445\;$ mm/s and $\alpha=0.531$, as shown in Fig. 5. The inset in Fig. 5 shows the quantity $(\Delta\gamma_{eff}/\Delta\gamma_{0}-1)$ in log-log scale to highlight the power law scaling of the effective surface energy with respect to the crack speed.
In summary, the dimensional parameters of the system are reported in Table 1. Notice that the natural frequency $\omega_{n}$ and the damping ratio $\zeta$ of the free oscillator (not in contact) were obtained performing a free vibration test of the indenter. In Table 1, the effective contact modulus $E^{*}_{eff}=4.5$ MPa is reported that is a factor $\simeq 1.6$ larger than the quasi-static value determined by the JKR contact test (see A). This accounts for the fact that the substrate gets stiffer when excited in the frequency range $f\in\left[200,400\right]$ Hz and agrees well with our vibrating experiments (see B) and with independent DMA characterization of the PDMS material in Maghami et al. (2024b).
@@
Gent-Shultz exponent $\alpha$ 0.531
@@ 4.4 Dynamic adhesion tests
Figure 6: Comparison with experiments a) Average spring reaction force ($\overline{F}_{k}$) vs. average contact radius ($\overline{a}$) curves for the dynamic tests performed with $f=300$ Hz and $X_{b}=[20,40,60,80,100]$ $\mu$m. b) Experimental pull-off force (3 test were performed per each value of amplitude $X_{b}$: data points correspond to the mean, the error bars indicate the standard deviation) vs. numerical prediction for different values of frequency ($f=[200,300,400]$ Hz). The black straight line indicates the locus where the experimental pull-off force $\overline{F}_{po}^{exp}$ is equal to the numerical pull-off force $\overline{F}_{po}^{num}$. c) Force ($\overline{F}_{k}$) vs. time (t) and contact radius (a) vs. time (t) curves for the case ($X_{b}=80$ $\mu$m, $f=300$ Hz). Experiments and simulations are matched at the time instant when unloading starts ($t\simeq 120$ s). Note that each marker shape always corresponds to the same value of $X_{b}$, in all the panels. The extra orange square and right-oriented triangle in panel b) refer to $X_{b}=120$ $\mu$m and $X_{b}=160$ $\mu$m, respectively
@@
In the classical JKR model, in load control, detachment happens through an instability at the turning point of the contact radius versus contact force curve where $dF_{c}/da$ vanishes. The experimental data at low vibration amplitude suggest the same mechanism as it can be seen from Fig. 6a. Increasing the vibration amplitude $X_{b}$ the detachment in the experiments appears more abrupt, suggesting a different mechanism.
@@
The behavior at pull-off was further scrutinized in Fig. 7a, where the pull-off force is shown as a function of the vibration amplitude for $f=300$ Hz. The black solid line refers to the model predictions while the markers (each corresponding to a certain value of $X_{b}$, consistently with Fig. 6) refer to the experimental data (the error-bar indicates one standard deviation over three independent measurements). Figure 7 b-i show micrographs of the contact area immediately prior to detachment, with the contact periphery indicated by a dashed orange circle. Pull-off is preceded by a significant reduction of contact radius in the quasi-static case and in the small amplitude cases. As $X_{b}$ increases, the detachment takes place in a more ”abrupt” manner. The latter case, referring to $X_{b}=[100,120,160]$ $\mu$m, seems related to the saturation of the pull-off force that the model is unable to predict. It appears that above a certain vibration amplitude, $X_{b}\simeq 80$ $\mu$m for the case considered in Fig. 7a, the condition from detachment shifts from an instability point to a different mechanism. Further scrutiny of the micrographs at high vibration amplitudes reveals the presence of radial wrinkles (see close-up view for the case $X_{b}=160$ $\mu$m in Fig. 7j) close to the contact periphery. A possible interpretation is that these wrinkles are the result of local elastic instabilities on the surface due to the hoop compressive stress $\sigma_{\theta}$ generated during the unloading phase, which for $\nu=0.5$, are equal to the Hertzian component of the pressure distribution acting normal to the substrate interface ($\sigma_{\theta}=-p_{0}\sqrt{1-(r/a)^{2}}$, being $p_{0}$ the maximum stress in the Hertzian solution, $r$ the radial coordinate, see Eq. (4.130) in Maugis (2013)). Nevertheless, local instabilities, resulting in wrinkles and cavitation phenomena, have been studied theoretically and numerically in spherical adhesive contact. He et al. (2022) observed them for quasi-static loading of very confined elastic systems where $a/h\gtrsim 4$, being $h$ the thickness of the layer (while in our case we have $a/h\simeq 1)$. However, our experimental set-up involves a viscoelastic material with base excitation, therefore definitive conclusions on the origin of the observed wrinkles cannot be drawn at this point.
@@
By considering both the pull-off force and the contact area (for $X_{b}>80$ $\mu$m, we considered only the patch within the white dashed circle to be in intimate contact), we estimated the average stress at pull-off, $\overline{\sigma}_{po}$, which is plotted as a function of the vibration amplitude in the inset of Fig. 7a. The results show that $\overline{\sigma}_{po}$ increases with $X_{b}$, peaks at $X_{b}=80$ $\mu$m, and decreases as $X_{b}$ increases further. This has two implications: firstly, there is an ”optimal” vibration amplitude that maximizes the adhesive capabilities of the interface; secondly, the pull-off force saturation cannot be ascribed to the limited adhesive strength of the interface, as otherwise we would have observed a plateau of $\overline{\sigma}_{po}$ at high vibration amplitudes (a ”strength-limited” mechanism, see Papangelo and Ciavarella (2023)). This, together with the observation of radial wrinkles, raises the question of what is the mechanism that triggers the saturation of the pull-off force at large amplitudes. This matter goes beyond the scope of the present work, and will be subject of further investigations.
Furthermore, the radial asymmetry of the wrinkles (indicated by the non-concentricity of the orange and white dashed circles in Fig. 7f-j) might suggest the presence of shear stresses. Oliver et al. (2023) showed that shear forces can reduce the pull-off force, nevertheless in our set-up the indenter is hung through a rubber band to the load cell like a pendulum, hence shear forces could not be equilibrated. Furthermore, shear forces introduce asymmetry into the contact area that appears more elliptical than circular (Sahli et al., 2019; Papangelo et al., 2019), something that we never observed in our experiments. Although a small misalignment in the set-up cannot be excluded, we believe the saturation of the pull-off force cannot be explained by the presence of shearing loads.
@@
Figure 7: a) Pull-off force as a function of the vibration amplitude $X_{b}$ (each marker corresponds to a certain value of $X_{b}$, consistently with Fig. 6). The model prediction (solid line) is compared against the experimental data points. b-i) Micrographs of the contact area just prior to pull-off for each test: b) quasi-static; c) $X_{b}=20$ $\mu$m; d) $X_{b}=40$ $\mu$m; e)$X_{b}=60$ $\mu$m; f) $X_{b}=80$ $\mu$m; g) $X_{b}=100$ $\mu$m; h) $X_{b}=120$ $\mu$m; i) $X_{b}=160$ $\mu$m; j) close-up view of the radial wrinkles leading to abrupt pull-off ($X_{b}=160$ $\mu$m). The frequency was set to 300 Hz. The contrast was enhanced to allow for a better visualization. In panels (b-j) the the orange dashed circles delimitate the contact periphery, while in panels (f-j) the white dashed circles indicate what we considered the intimate contact circle, i.e. not including the radial wrinkles.
@@ 5 Discussion
In contrast, here we started with an independent characterization of the rate-dependent adhesion law under quasi-static conditions (see Fig. 5), i.e., without vibrations. By providing all the dimensional characteristic parameters of our experimental setup as input to the model, it successfully predicted the experimental results, including the time-dependent evolution of the contact radius, spring reaction force and the pull-off force, without any adjustable parameter. Nevertheless, in experiments, the pull-off force saturated above an excitation amplitude of about $100$ $\mu$m. This saturation, which could not be predicted by the model, may be related to the onset of local surface instabilities detected in the contact area snapshot at the pull-off point (see Fig. 7) perhaps due to local circumferential compressive stresses arising within the contact area. Notice that, once the indenter detached from the substrate, we could not identify any permanent damage on the PDMS surface. In order to explain the saturation of the pull-off force, Shui et al. (2020) have proposed a truncation of the Gent and Schultz rate-dependent adhesion law at a critical crack speed $v_{C}$. This scenario does not find an experimental evidence in our measurements, where the effective work of adhesion was found to monotonically increase with the crack speed (see Fig. 5). Moreover, the pull-off saturation we found appeared to be frequency dependent (see Fig. 6b), which does not support the existence of a speed cut-off in the Gent and Schultz law, which would have appeared as an inherent property of the interface providing the same pull-off saturation independently on the excitation frequency.”
@@
Notice that, we have found that the system response is apparatus-dependent in the range of excitation frequency close to the system resonance, while if $f>>f_{nc}$ there is no dependence on the indenter mass and apparatus stiffness as we discuss in Appendix C.
@@ 6 Conclusions
To quantitatively compare with the model predictions, we conducted an independent assessment of the rate-dependent adhesion performance of the glass sphere-PDMS substrate interface. The rigid sphere was unloaded from the PDMS substrate at several unloading rates, and the relation between the effective surface energy, $\Delta\gamma_{eff}$, and the crack speed was obtained. During the characterization, the vibration excitation of the substrate was switched off. Having obtained the parameters of the Gent and Schultz adhesion law ($v_{0}=44.5$ $\mu$m/s and $\alpha=0.531$) for the interface, we exploited the mechanical model to predict the experimental results. This resulted in a very satisfactory comparison, not only for single-value predictions such as the pull-off force, but also for the entire evolution of average contact radius and contact force over time.
@@ Acknowledgment
All authors were partly supported by the Italian Ministry of University and Research under the Programme “Department of Excellence” Legge 232/2016 (Grant No. CUP - D93C23000100001). A.P. and M.T. were supported by the European Union (ERC-2021-STG, “Towards Future Interfaces With Tuneable Adhesion By Dynamic Excitation” - SURFACE, Project ID: 101039198, CUP: D95F22000430006). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.C. was partly supported by the European Union through the program – Next Generation EU (PRIN-2022-PNRR, ”Fighting blindness with two photon polymerization of wet adhesive, biomimetic scaffolds for neurosensory REtina-retinal Pigment epitheliAl Interface Regeneration” - REPAIR, Project ID: P2022TTZZF, CUP: D53D23018570001). A.P. is thankful to Julien Scheibert (LTDS, CNRS, Lyon) for insightful discussions and for the support received to set-up the experimental test-rig.
@@ Appendix A Characterisation of the PDMS sample
Figure 8: Average spring reaction force ($\overline{F}_{k}$) vs. average contact radius ($\overline{a}$) curves for the quasi-static test (gray dots) and dynamic experimental tests (markers) performed at $f=300$ Hz and $X_{b}=[20,40,60,80,100]$ $\mu$m with the corresponding JKR fits (black dashed lines). Notice that the black dashed lines reported should be interpreted as JKR curves with effective properties, while the solid lines in Fig. 6a were obtained upon integration of the full dynamical model (Eq. (16)).
Figure 9: Evolution of the effective contact modulus $E^{*}_{eff}$ and effective work of adhesion $\Delta\gamma_{eff}$ as functions of the vibration amplitude $X_{b}$ ($f=300$ Hz). The data points were obtained by fitting the experimental force vs. contact radius curves at unloading with a JKR model, as shown in Fig. 8
Appendix B Comparison with Shui et al. (2020)
Shui et al. (2020) has proposed a slightly different model for the system in Fig. 1 starting from the hypothesis that the effect of the vibrating substrate can be taken into account by a forcing term $f(t)\simeq-m\omega^{2}X_{b}\sin\left(\omega t\right)$. If, in the model proposed here, we write the indentation as a function of the oscillator and base position
@@
From the JKR model writing the contact force $F_{c}=\frac{Ka}{2}\left(\frac{a^{2}}{R}-3\delta\right)$, with $K=4/3E^{*}$, and considering $k\rightarrow 0$ and $c\ll m\omega$, one obtains
@@
Nevertheless, the predictions of Eq. (25) compared very poorly with our experimental data (see Fig. 10). The striking difference between the model proposed here and the approximate solution derived by Shui et al. (2020) is in the definition of the damping coefficient $c$. In the model proposed here $c$ includes the dissipative phenomena at play when the indenter is freely oscillating (not in contact), while the dissipative contribution coming from the viscoelastic bulk have been independently characterized by the Gent and Schultz law. Conversely, the model proposed by Shui et al. (2020), incorporates within the damping coefficient $c$ all the dissipative phenomena of the system, including the hysteresis originated from the viscoelastic nature of the bulk material. In the supplementary information of their paper, Shui et al. (2020) suggest that the damping coefficient should be determined from the input power provided to the system, nevertheless it remains unclear how this should be related to the rate-dependent adhesive behavior of the soft interface.
Figure 10: Pull-off force as a function of the vibration amplitude $X_{b}$. The markers represent the experimental data, the present model is shown as a black solid line while the predictions of Eq. (25) by Shui et al. (2020) is shown as an orange solid line. The case shown corresponds to $f=300$ Hz.
@@
being $\omega_{nc}=\sqrt{\left(\left.\frac{dF_{c}}{d\delta}\right|_{\delta=\overline{\delta}}\right)/m}$ which gives the ”resonance” frequency of the system. Notice that the latter is $\overline{\delta}$ dependent. Equation (26) is only a simple linearized approximation of the system dynamical behavior. A full treatment of the nonlinear dynamical system response is out of the scope of the present work.